What I need is a thorough laymens term on how to solve this math.
Determine whether the point (7,2) is a solution of the system
-3x+4y=-13
-2x+8y=2
by substituting the point into each equation?
Thank you so much

Answer by stanbon(75887)

What I need is a thorough laymens term on how to solve this math.
Determine whether the point (7,2) is a solution of the system
-3x+4y=-13
-2x+8y=2
by substituting the point into each equation?
---------------
Yes. If the point is a solution for the system it
must lie at the intersection of the two lines.
(7,2) must make both equations "true".
----
See if it does.
Substituting in each equation you would get:
-3*7+4*2 = -13
-21+8 = -13
-13 = -13
OK
------------------------
-2*7+8*2 = 2
-14 + 16 = 2
2 = 2
OK
------------------------
It makes both equations "true" so it is the
solution for the system.
==================================
cheers,
Stan H.